Mr. Sam Golff desires to invest a portion of his assets in rental property. He has narrowed his choices down to two apartment complexes, Palmer Heights and Crenshaw Village. After conferring with the present owners, Mr. Golff has developed the following estimates of the cash flows for these properties. Palmer Heights Yearly Aftertax Cash Inflow (in thousands) Probability $ 140 .1 145 .2 160 .4 175 .2 180 .1 Crenshaw Village Yearly Aftertax Cash Inflow (in thousands) Probability $ 145 .2 150 .3 160 .4 170 .1 a. Find the expected cash flow from each apartment complex. (Enter your answers in thousands (e.g, $10,000 should be enter as "10").) b. What is the coefficient of variation for each apartment complex? (Do not round intermediate calculations. Round your answers to 3 decimal places.) c. Which apartment complex has more risk? Palmer Heights Crenshaw Village
Answer #1

1.
Palmer Heights=0.1*140+0.2*145+0.4*160+0.2*175+0.1*180=160

Crenshaw Village=0.2*145+0.3*150+0.4*160+0.1*170=155

2.
Palmer Heights=sqrt(0.1*(140-160)^2+0.2*(145-160)^2+0.4*(160-160)^2+0.2*(175-160)^2+0.1*(180-160)^2)/160=0.08149003

Crenshaw Village=sqrt(0.2*(145-155)^2+0.3*(150-155)^2+0.4*(160-155)^2+0.1*(170-155)^2)/155=0.049973979

3.
Palmer Heights